OpenStudy (anonymous):

3x - 2y = 10 5x + y = 4 When solving this system of equations by elimination, which could be the resulting equation when a variable has been eliminated?

OpenStudy (anonymous):

Can you elaborate? Or perhaps include the original question(s)?

OpenStudy (anonymous):

this is the original question :/ do you awnt to see the answer choices?

OpenStudy (anonymous):

Is this from a textbook? Or a review sheet?

OpenStudy (anonymous):

an online thing

OpenStudy (anonymous):

Can you show me the answer choices? I've got an answer or two, but they seem funky.

OpenStudy (anonymous):

yep! one second

OpenStudy (anonymous):

A) 13x = 18 B) -7x = 2 C) -7y = 62 D) 8x - y = 14

OpenStudy (anonymous):

Ok. I got "A."

OpenStudy (anonymous):

Let me show you my steps

OpenStudy (anonymous):

sweet. thank you so much!

OpenStudy (anonymous):

all right.

OpenStudy (anonymous):

\[1.) (3x-2y=10) x 1 --> I multiplied by 1, to leave the equation the same (5x+y=4) x 2 --> I multiplied by 2 in order to make the "y" a "-2y"

OpenStudy (anonymous):

You now have: 3x-2y=10 10x+2y=4 --> Distributed the 2 into the whole equation. (notice the "-2" mentioned in the above step.

OpenStudy (anonymous):

2.) Add the two equations. (3x) + (10x) = 13x (2y) + (-2y) = 0 (This is why we multiplied by 2.) (8)+(10)= 18 You'll end up with: 13x=18 Hope this helped! :D

OpenStudy (anonymous):

yeah it helped a bunch! thanks a lot